A spring of a natural length 3cm is extended by 100cm by a force of 9N,what will be its length when the applied force is10N.


1
Expert's answer
2021-10-12T12:31:48-0400

Given:

"l_0=0.03\:\rm m"

"l_1=1.03\:\rm m"

"F_1=9\:\rm N"

"F_2=10\:\rm N"


The Hooke's law says

"F=k(l-l_0)"

Hence


"\frac{F_2}{F_1}=\frac{l_2-l_0}{l_1-l_0}"

"l_2=l_0+\frac{F_2}{F_1}(l_1-l_0)\\
=0.03+\frac{10}{9}*1.0=1.14\:\rm m=114\: cm"